OpenStudy (anonymous):

Which is an equation in standard form of a line that passes through (2, –1) and has a slope of

OpenStudy (anonymous):

to solve this I would need the slope, but you want to use point slope form which is \[y-y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

where m= slope, y1=y coordinate, x1= x coordinate

OpenStudy (anonymous):

you plug in the information you are given and solve for y, not y1, y

OpenStudy (anonymous):

y-1=m(x-2) I believe

OpenStudy (anonymous):

ya, from the info we've been given so far @TheTerrence is what we have but to solve this and put it in standard form we need the slope

OpenStudy (anonymous):

So Ax+By=C then?

OpenStudy (anonymous):

@TechnoSoul

OpenStudy (anonymous):

i think they just forgot to put the slope i mean in this way of wording the problem it seems like the slope was given